OpenStudy (anonymous):

With a certain tail wind a jet aircraft arrives at its destination, 1,890 miles away, in 3 hours. Flying against the same wind, the plane makes the return trip in 3 3/8 hours. Find the wind speed and the plane's airspeed. {wind speed is a0 mph, airspeed is a1 mph}

OpenStudy (anonymous):

put the airspeed as \(a_1\) and the wind speed as \(a_0\) then the speed going is \[a_1+a_0\] and the speed returning is \(a_1-a_0\)

OpenStudy (anonymous):

since distance is rate time time, you have \[1890=3(a_1+a_0)\] and also \[1890=3.85(a_1-a_0)\] solve the two equations in two unknowns